OpenStudy (anonymous):

The value of y varies directly with x, and y = 4 when x = 1.Find y when x = −3.

OpenStudy (shamil98):

First you need to find k. k = y/x k = 4/1 k = 4 y = kx y = 4(-3) y = -12

OpenStudy (shamil98):

Take a look at http://www.regentsprep.org/Regents/math/ALGEBRA/AO4/Ldirect.htm for more information on direct variation..